OpenStudy (fanduekisses):

How do I Find lim x->o of Sinx/5x ?

OpenStudy (anonymous):

L'opital

OpenStudy (anonymous):

L'hopital

OpenStudy (fanduekisses):

\[\lim_{x \rightarrow 0} \frac{ sinx }{ 5x}\]

OpenStudy (fanduekisses):

What is it? how does it work?

OpenStudy (anonymous):

Search it

OpenStudy (fanduekisses):

ok

OpenStudy (anonymous):

It's because I'm not good at english xD

OpenStudy (anonymous):

take the \(\frac{1}{5}\) outside the limit as it is a constant

OpenStudy (anonymous):

\[\frac{1}{5}\overbrace{\lim_{x\to 0}\frac{\sin(x)}{x}}^{\text {you are supposed to know this one}}=\frac{1}{5}\times 1=1\]

OpenStudy (fanduekisses):

oooohhh I understand now

OpenStudy (fanduekisses):

thanks

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

except i made a dumb mistake, the answer is \(\frac{1}{5}\) not \(1\)

OpenStudy (fanduekisses):

yeah I caught it :)